if the samples are 16, 17, 15, 24, 20, 20, 20, 23, 18, and 19. what is the point estimate of the population standard deviation?




1
Expert's answer
2022-04-21T10:07:38-0400

The sample standard deviation s is a point estimate of the population standard deviation σ:

"s=\sqrt{\cfrac{\sum{( x-\bar x)^2}}{n-1}};\\
\bar x=\\
=\cfrac{16+17+15+24+20+20+20+23+18+19}{10}=\\
=19.2;\\"

"X-\bar x=\\
=\begin{Bmatrix}
 16-19.2, 17-19.2,15-19.2,...,19-19.2
\end{Bmatrix}=\\
=\begin{Bmatrix}-3.2,-2.2,-4.2,...,-0.2\end{Bmatrix};"


"\sum{( x-\bar x)^2}=(-3.2)^2+(-2.2)^2+(-4.2)^2+...+\\
+(-0.2)^2=73.6;\\
s=\sqrt{\cfrac{73.6}{10-1}}=2.86."